OpenStudy (anonymous):

What is the monthly payment of a financed amount of $9,750 with an APR of 8.9 percent? The term is for 3 years and the monthly payment per $100 is $3.52. A. $2,605.20 b.$348.29 c.$12,355.20 d. $241.00

OpenStudy (amistre64):

ideally, you can work a recurrsion into the explicit result of:\[B_n=B_o(k)^n-P\frac{1-k^n}{1-k}~:~k=1+\frac{r}{m}~:~n=m*years\]

OpenStudy (amistre64):

sposing that the final payment is to zero out the balance ... \[0=B_o(k)^{36}-P\frac{1-k^{36}}{1-k}\] \[B_o(k)^{36}=P\frac{1-k^{36}}{1-k}\] \[B_o(k)^{36}\frac{1-k}{1-k^{36}}=P\] \[9750(1+\frac{.089}{12})^{36}\frac{1-(1+\frac{.089}{12})}{1-(1+\frac{.089}{12}) ^{36}}=P\]

OpenStudy (amistre64):

ideally of course, but then i dont know the method they expect you to employ on this. with any luck its close to 310